OpenStudy (anonymous):

Find the value of x. A. 7 B. 9 C. 11 D. 12

OpenStudy (anonymous):

Ok

OpenStudy (unklerhaukus):

The angle bisector theorem says: "If ABC is any triangle, and BD bisects the angle ABC, then AD/CD = AB/BC"

OpenStudy (unklerhaukus):

\[\frac x{16-x}=\frac{36}{28}\] solve for x